OpenStudy (anonymous):

The sum of squares formula is given by 1^2+2^2+3^2+…+n^2=n(n+1)(2n+1)/6. The sum of odd squares can be expressed as 1^2+3^2+5^2+…+(2n−1)^2=An^3+Bn^2+Cn+D. The value of A can be expressed as a/b, where a and b are positive coprime integers. What is the value of a+b?

OpenStudy (raden):

1^2+3^2+5^2+…+(2n−1)^2 = 4/3 n^3 - 1/3 n therefore, A = 4/3 = a/b so, a+b = .... ?

OpenStudy (anonymous):

is it 5?????

OpenStudy (raden):

a/b = 4/3 obviously a=4 and b=3 (with a and b are positive coprime integers) so, a+b = 4 + 3 = ... ?

OpenStudy (anonymous):

7

OpenStudy (raden):

yes